OpenStudy (anonymous):

please solve..Find the midpoint of the segment with the given endpoints. M(–2, 1) and O(–3, 2)

OpenStudy (anonymous):

thank youuu can you help me with another problem?

OpenStudy (anonymous):

find the midpoint of line ab

OpenStudy (anonymous):

Actually, Midpoint = total/2

OpenStudy (anonymous):

? huh

OpenStudy (anonymous):

=> Midpoint ( -5/2 , 3/2 )

OpenStudy (anonymous):

not an option these are A. (3, 1) B. (1, 1) C. (10, 4) D. (4, 4)

OpenStudy (anonymous):

@kde123 sorry, @Chlorophyll is right. I forgot :) The formula for mid point is mid x= (x1+x2)/2 mid y= (y1+y2)/2

OpenStudy (anonymous):

the answer is?

OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. M(–2, 1) and O(–3, 2) => Midpoint ( -5/2 , 3/2 )

OpenStudy (anonymous):

@kde123 One question per post, keep it less confusing!

OpenStudy (anonymous):

okayy im sorry im confused now i just want the mid point of ab the graph is attached

OpenStudy (anonymous):

A. (3, 1) B. (1, 1) C. (10, 4) D. (4, 4)